3. In a dissolution experiment, 5.0 g of NH4NOs was dissolved in 50.0 mL of water in a calorimeter If the calculated for the reaction was 1.7 kJ, calculate the observed Ar for the experiment. (Ignore the heat capacity of the calorimeter, i.e., use Ceup, -0. Assume the specific heat capacity of the solution = 4.18 J/g。C and the density of the water-1.0 g/mL)
Answer #1

qdissolution= -Qmix

Qmix = -106.182 J

Q = m*C*(Tf-Ti)

-1.7*10^3= 50*4.184*(dT)

dT = -1700/( 50*4.184)

dT = -8.126°C